# Solvable families of random block tridiagonal matrices

### v1

@@ 1 Introduction
Let $\mathtt{H}_{\beta,n}(r,s)$ denote the distribution of the $rn\times rn$ symmetric or Hermitian block tridiagonal matrix with $r\times r$ diagonal blocks distributed as independent copies of G(O/U)E, and descending upper diagonal blocks distributed as independent copies of the (lower triangular) positive square root of a real/complex Wishart with parameters $(r,(r+s)(n-i))$. Here $i$ is the index of the offdiagonal block entry, and $\beta=1$ and 2 corresponds to the real and complex case, respectively. As in the $r=1$ case, the diagonal and and offdiagonal variables are also independent of each other. A more detailed description of these ensembles is provided in Section 2.2.
@@
Here for $r\geq 2$ and $n\geq 1$, ${\mathcal{P}}_{r,n}$ denotes the set of size $r$ equipartitions of $[\![rn]\!]:=\{1,2,\dots rn\}$. That is, $\{\mathcal{A}_{1},\dots\mathcal{A}_{r}\}\in{\mathcal{P}}_{r,n}$ if $|\mathcal{A}_{i}|=n$ for all $i$ and the $\mathcal{A}_{i}$ form a partition of $[\![rn]\!]$. With that, for any $\mathcal{A}\subset[\![rn]\!]$, we write $\Delta(\mathcal{A})$ as shorthand for the Vandermonde determinant in the $|\mathcal{A}|$ ordered eigenvalue variables with indices drawn from $\mathcal{A}$ (suppressing the explicit dependence on $\lambda_{i},i\in\mathcal{A}$). Finally, $\operatorname{Pf}(M)$ denotes the Pfaffian of $M$.
@@ 1.1 Soft edge asymptotics of $\mathtt{H}_{\beta,n}(r,s)$
When $\gamma=1$ this is the multivariate Stochastic Airy Operator of [2]. In particular, with a Dirichlet boundary condition at the origin, the spectrum of $-\mathcal{H}_{\beta}=-\mathcal{H}_{\beta,1}$ is given by the $\operatorname{Airy}_{\beta}$ process, the edge scaling limit of the Gaussian beta ensemble. The largest value of this process (which is minus the ground state eigenvalue of $\mathcal{H}_{\beta}$), has classical Tracy-Widom distribution $TW_{\beta}$ with $\beta=1,2,4$.
@@
and by $\Lambda_{0}<\Lambda_{1}<\cdots$ the Dirichlet eigenvalues of $\mathcal{H}_{\beta,\gamma}$ with the choice $\gamma=\frac{r+s}{r}$ Then the point process $\{\lambda_{0}^{(n)},\lambda_{1}^{(n)},\dots\}$ converges in law to $\{\Lambda_{0},\Lambda_{1},\dots\}$ as $n\to\infty$.
@@
starting from $(p_{1}(0),\cdots,p_{r}(0))=\{\infty\}^{r}$ and entering $\{p_{1}<\cdots<p_{r}\}$ at $x>0$. Here $(b_{1},\cdots b_{k})$ is a standard real $r$-dimensional Brownian motion; $p_{1}$ can hit $\infty$- ∞ in finite time, whereupon it is placed at +$\infty$+ ∞ and the re-indexed process starts afresh. Then with $\Lambda_{0}<\Lambda_{1}<\cdots$ defined as in Theorem 2, it holds that
@@ 1.2 Block Laguerre ensembles
When $\beta=1$ or $2$ this coincides with that of the law of a sample covariance matrix for $m\geq n$ independent real or complex normal samples in dimension $n$. Along with $\beta$ now taking any positive value, the model behind (7) allows $m$ to be generalized to any real number greater than $n-1$.
We define the distribution $\mathtt{W}_{n,m,\beta}(r,s)$ on nonnegative definite block tridiagonals as follows. Let $\mathbf{L}_{n}$ be an $rn\times rn$ block bidiagonal matrix with independent $r\times r$ diagonal and upper offdiagonal blocks denoted by $\{\mathbf{D}_{i}\}_{i=1,n}$ and $\{\mathbf{O}_{i}\}_{i=1,n-1}$, that are lower and upper triangular matrices, respectively. Distribute these according to square-root Wishart matrices with parameters $(r,(r+s)(m+1-i))$ and $(r,(r+s)(n-i))$, respectively. Then $\mathtt{W}_{n,,m,\beta}(r,s)$ has the law $\mathbf{L}_{n}\mathbf{L}_{n}^{\dagger}$. Full details are provided in Definition 20.
@@
The joint eigenvalue density of $\mathtt{W}_{n,m,\beta}(r,s)$ for $\beta=1$ or $2$ has the form (2) for general $r\geq 2$ and $\beta s=2$ and (3) for $r=2$ and $\beta s=2$ or $4$ with an explicitly computable normalizing constant, the only change being that the Gaussian weight $e^{-\frac{\beta}{4}\sum_{i=1}^{rn}\lambda_{i}^{2}}$ is replaced by $\prod_{i=1}^{rn}\lambda_{i}^{\frac{\beta}{2}((r+s)(m-n)+1)-1}e^{-\frac{\beta}{2}\lambda_{i}}$, restricted to ${\mathbb{R}}_{+}^{rn}$.
In terms of asymptotics, we focus on the choice $m=n+a$ for fixed $a>-1/(r+s)$ as $n\rightarrow\infty$ and look at the scaling limit of the smallest eigenvalues, which end up being in the vicinity of the origin. This is the random matrix hard edge, and introduces novel limiting phenomena beyond what we have seen for $\mathtt{H}_{n,\beta}(r,s)$. Note that it may proved along the same lines to Theorem 2 that the suitably centered and scaled largest eigenvalues under $\mathtt{W}_{n,m,\beta}(r,s)$ will converge to those of $\mathcal{H}_{\beta,\gamma}$, for an appropriate $\gamma$, and the same is in fact true for the smallest eigenvalues when $\liminf_{n\to\infty}m/n>1$.
@@
Let $\mathbb{P}$ be the measure on (non-intersecting) paths ${\mathbf{q}}:[\mu,\infty)\mapsto[-\infty,\infty]^{r}$ defined by
$dq_{i}=\frac{2}{\sqrt{\beta\gamma}}q_{i}db_{i}+\left(\left(\frac{a}{\gamma}+\frac{2}{\beta\gamma}\right)q_{i}-q_{i}^{2}-e^{-rx}+q_{i}\sum_{j\neq i}\frac{q_{i}+q_{j}}{q_{i}-q_{j}}\right)dx,$ (9)
started from {}$\{\infty\}^{r}$ with the same ordering and re-indexing conventions upon possible passages to $\infty$- ∞ described in Corollary 3. With $0<\hat{\Lambda}_{0}<\hat{\Lambda}_{1}<\cdots$ defined in Theorem 7 it holds
$P(\hat{\Lambda}_{k}>\lambda)=\mathbb{P}(x\mapsto\mathbf{q}(x)\mbox{ vanishes at most }k\mbox{ times })$ (10)
@@ 2.1 Block Jacobi matrices
[…]
@@ 2.2 Random block matrices
Let $r,n\geq 1$, $m>-1/r$, and $s\geq 0$. Let $\mathbf{L}$ be an $rn\times rn$ block bidiagonal matrix with independent $r\times r$ diagonal and upper offdiagonal blocks denoted by $\{\mathbf{D}_{i}\}_{i=1,n}$ and $\{\mathbf{O}_{i}\}_{i=1,n-1}$ with $\mathbf{D}_{i}^{{\dagger}}\sim\mathtt{SqW}_{\beta}(r,(r+s)(m+1-i))$ and $\mathbf{O}_{i}\sim\mathtt{SqW}_{\beta}(r,(r+s)(n-i))$. We denote the distribution of $\mathbf{W}=\mathbf{L}\mathbf{L}^{{\dagger}}$ by $\mathtt{W}_{\beta,n,m}(r,s)$.
@@ 3.1 Revisiting the Hermite beta ensemble
Moreover, the spectral weights of $\mathbf{T}$ corresponding to the first coordinate vector have Dirichlet$(\beta/2,\dots,\beta/2)$ joint distribution, and this weight vector is independent of the eigenvalues.
@@
hence biasing the entries of $\mathbf{T}$ with $g_{\beta}(\mathbf{b})$ is the same as biasing the spectral variables $\lambda,\mathbf{p}$ with the appropriate product on the right hand side of (26). This immediately implies that the eigenvalues and spectral weights of $\mathbf{\widetilde{T}}$ are still independent of each other, that the joint eigenvalue density of $\mathbf{\widetilde{T}}$ is proportional to $|\Delta(\lambda)|^{\beta}e^{-\frac{1}{4}\sum_{k=1}^{n}\lambda_{k}^{2}}$, and that its spectral weights have Dirichlet$(\beta/2,\dots,\beta/2)$ distribution.
@@ 3.2 Key spectral identity
Hence $\mathbf{S}={\mathbf{Q}}\widetilde{\mathbf{S}}$, and $|\det\mathbf{S}|=|\det\widetilde{\mathbf{S}}|$ with
@@ 3.3 Proofs of Theorem 1 and 6
×$\times\prod_{j=1}^{n-1}(\det\mathbf{O}_{j})^{\beta(r+s)(n-j)+\beta-1}\prod_{j=1}^{n}\prod_{k=1}^{r}(d^{(j)}_{k,k})^{-\beta k}\prod_{j=1}^{n-1}\prod_{k=1}^{r}(o^{(j)}_{k,k})^{-\beta k},$
@@ 3.4 A reduction and a representation
For $\underline{\mathcal{A}}=(\mathcal{A}_{1},\dots,\mathcal{A}_{r})\in{\mathcal{P}}_{r,n}$ let $\sigma_{\underline{\mathcal{A}}}\in S_{rn}$ denote the following permutation of the set $[\![rn]\!]$:
@@
Now when $r=2$ the elements of $(\mathcal{A}_{1},\mathcal{A}_{2})$ of ${\mathcal{P}}_{2,n}$ can be identified just with the subsets $\mathcal{A}\subset[\![2n]\!]$ of cardinality $n$. Using $\mathcal{A}^{\prime}$ for the complement of $\mathcal{A}$ in $[\![2n]\!]$ throughout, we further denote:
@@
First, any permutation $\sigma\in S_{rn}$ of $[\![rn]\!]$ can be mapped in a one-to-one way to a pair consisting of an element of ${\mathcal{P}}_{r,n}$ and a vector $(\sigma_{1},\dots,\sigma_{r})$ of $r$ permutations of $S_{n}$. For $\sigma\in S_{rn},m\in[\![r]\!]$ let
@@
×$\hskip 130.0pt\times\Delta(\mathcal{A})\Delta(\mathcal{A}^{\prime})\Delta(\mathcal{B})\Delta(\mathcal{B}^{\prime}).$
@@
×$\hskip 80.0pt\times\sum_{a=0}^{n}g(a,a,2n-a)\sum_{|\mathcal{B}|=n,|\mathcal{A}\cap\mathcal{B}|=a}\operatorname{sgn}(\sigma_{\mathcal{B}})\Delta(\mathcal{B})\Delta(\mathcal{B}^{\prime}).$
@@
To prove (60) in the $11\leq a\leq n-$ 1 case we use the following identity due to Sylvester [24]. Suppose that $\mathbf{R}_{1}$, $\mathbf{R}_{2}$ are $n\times n$ matrices and $1\leq m\leq n$. Let $I$ be a fixed subset of column indices of $\mathbf{R}_{1}$ with $|I|=m$. Then
@@
Fix $11\leq a\leq n-$ 1, $\mathcal{A}\subset[\![2n]\!]$ with $|\mathcal{A}|=n$, and $\mathcal{S}\subset\mathcal{A}$ with $|\mathcal{S}|=a$. Let $\mathbf{R}_{1}$ be the $n\times n$ Vandermonde matrix corresponding to $\mathcal{A}$, $\mathbf{R}_{2}$ the Vandermonde matrix corresponding to $\mathcal{A}^{\prime}$, and consider Sylvester’s identity applied for $m=n-a$ where the fixed set of coordinates of $\mathbf{R}_{1}$ correspond to $\widetilde{\mathcal{S}}:=\mathcal{A}\setminus\mathcal{S}$. For each $(\widehat{\mathbf{R}}_{1},\widehat{\mathbf{R}}_{2})$ in the sum, let $\mathcal{B}$ be the set of indices corresponding to the columns of $\widehat{\mathbf{R}}_{1}$, then $|\mathcal{B}|=n,\mathcal{A}\cap\mathcal{B}=\mathcal{S}$. We set $\mathcal{T}=\mathcal{A}^{\prime}\cap\mathcal{B}^{\prime}$ and $\widetilde{\mathcal{T}}=\mathcal{A}^{\prime}\setminus\mathcal{T}$. The left side of (62) is $\Delta(\mathcal{A})\Delta(\mathcal{A}^{\prime})$, while the sum on the right side is
@@
We will do this by transforming the permutation $\sigma_{\mathcal{A}}$ into $\sigma_{\mathcal{B}}$ while keeping track of the signature of the required steps. Consider the permutation $\sigma_{\mathcal{A}}$. Interchange the indices in $\widetilde{\mathcal{S}}$ with $\widetilde{\mathcal{T}}$ (keeping their respective order), since $|\widetilde{\mathcal{S}}|=|\widetilde{\mathcal{T}}|=n-a$, we can achieve this with $n-a$ transpositions. Then reorder the elements in positions $11,3,\dots,2n-$ 1, and then in the positions $2,4,\dots,2n$. We can do this using the permutations $\sigma_{\mathcal{A},\widetilde{\mathcal{S}}\to\widetilde{\mathcal{T}}}$ and $\sigma_{\mathcal{A}^{\prime},\widetilde{\mathcal{T}}\to\widetilde{\mathcal{S}}}$ applied to these collections of indices, so the signature of this step is exactly $\operatorname{sgn}(\sigma_{\mathcal{A},\widetilde{\mathcal{S}}\to\widetilde{\mathcal{T}}})\operatorname{sgn}(\sigma_{\mathcal{A}^{\prime},\widetilde{\mathcal{T}}\to\widetilde{\mathcal{S}}})$. The resulting permutation is exactly $\sigma_{\mathcal{B}}$, and the signature of the steps we have taken is $(-1)^{n-a}\operatorname{sgn}(\sigma_{\mathcal{A},\widetilde{\mathcal{S}}\to\widetilde{\mathcal{T}}})\operatorname{sgn}(\sigma_{\mathcal{A}^{\prime},\widetilde{\mathcal{T}}\to\widetilde{\mathcal{S}}})$, which proves (64). ∎
@@
$\mathcal{B}_{(j_{1},j_{2},j_{3},j_{4})}=\mathcal{A}_{1}^{*}\mathcal{A}_{2}^{*}\mathcal{A}_{3}^{*}\mathcal{A}_{4}^{*},\quad\text{with}\quad\mathcal{A}_{i}^{*}=\begin{cases}\mathcal{A}_{i},&\qquad\text{if }j_{i}=1,\\ \mathcal{A}_{i}^{\prime},&\qquad\text{if }j_{i}=0.\end{cases}$ (67)
@@
In other words, after we take expected value in (66), only those quadruples $(\mathcal{A}_{1},\dots,\mathcal{A}_{4})$ contribute for which intersections of the form $\mathcal{A}_{i_{1}}^{\prime}\mathcal{A}_{i_{2}}\mathcal{A}_{i_{3}}\mathcal{A}_{i_{4}}$ and $\mathcal{A}_{i_{1}}\mathcal{A}_{i_{2}}^{\prime}\mathcal{A}_{i_{3}}^{\prime}\mathcal{A}_{i_{4}}^{\prime}$ are empty (for distinct $i_{1},\dots,i_{4}$). Let $\mathcal{G}_{n}$ denote the set of quadruples $(\mathcal{A}_{1},\dots,\mathcal{A}_{4})$ that satisfy this compatibility condition together with $|\mathcal{A}_{i}|=n$, $\mathcal{A}_{i}\subset[\![2n]\!]$.
@@
$\begin{array}[]{c|c|c|c|c|c|c|c}1111&1100&1010&1001&0110&0101&0011&0000\\ \hline\cr 1&2&3&4&5&6&7&8\\ \end{array}$ (71)
Hence $\mathcal{B}_{(1,0,1,0)}$ and $\mathcal{B}_{3}$ both denote the set $\mathcal{A}_{1}\mathcal{A}_{2}^{\prime}\mathcal{A}_{3}\mathcal{A}_{4}^{\prime}$, and $b_{3}=b_{(1,0,1,0)}$. At the same time, if $(\mathcal{A}_{1},\dots,\mathcal{A}_{4})\in\mathcal{G}_{n}$, then each $\mathcal{A}_{i}$ and $\mathcal{A}_{i}^{\prime}$ can be written as the disjoint union of four of the sets $\mathcal{B}_{j},1\leq j\leq 8$. For example, $\mathcal{A}_{1}=\mathcal{B}_{1}\cup\mathcal{B}_{2}\cup\mathcal{B}_{3}\cup\mathcal{B}_{4}$. From $|\mathcal{A}_{i}|=|\mathcal{A}_{i}^{\prime}|$, $1\leq i\leq 4$, we get four equations with the sum of four terms on both sides. For instance, $|\mathcal{A}_{1}|=|\mathcal{A}_{1}^{\prime}|$ yields
@@
Next, for a particular $(\mathcal{A}_{1},\dots,\mathcal{A}_{4})\in\mathcal{G}_{n}$ consider the following “special” ordering of the elements of $[\![2n]\!]$: $x\prec y$ if $x\in\mathcal{B}_{i},y\in\mathcal{B}_{j}$ with $i<j$ or if $x<y$ with $x,y\in\mathcal{B}_{i}$. Denote by $\widetilde{\Delta}(\mathcal{A})$ the Vandermonde determinant corresponding to the elements $\lambda_{i},i\in\mathcal{A}$ using our special ordering, and let $\sigma_{i}$ be the permutation that maps $11,3,\dots,2n-$ 1 and $2,4,\dots,2n$ into the elements of $\mathcal{A}_{i}$ and $\mathcal{A}_{i}^{\prime}$ according to the special order. These permutations can be visualized using the following table:
@@
Here $a_{1},a_{2}$ are indices from the “$\mathcal{B}$-partition” of $\mathcal{A}_{i}$ and $b_{1},b_{2}$ are indices from the the “$\mathcal{B}$-partition” of $\mathcal{A}_{i}^{\prime}$. Multiplying these equations for $1\leq i\leq 4$ and some bookkeeping yields

### v2

@@ 1 Introduction
Let $\mathtt{H}_{\beta,n}(r,s)$ denote the distribution of the $rn\times rn$ symmetric or Hermitian block tridiagonal matrix with $r\times r$ diagonal blocks distributed as independent copies of G(O/U)E, and descending upper diagonal blocks distributed as independent copies of the (lower triangular) positive square root of a real/complex Wishart with parameters (r,(r+s)(ni))(r,(r+s)(n-i)). Here $i$ is the index of the offdiagonal block entry, and $\beta=1$ and 2 corresponds to the real and complex case, respectively. As in the $r=1$ case, the diagonal and and offdiagonal variables are also independent of each other. A more detailed description of these ensembles is provided in Section 2.2.
@@
Here for $r\geq 2$ and $n\geq 1$, ${\mathcal{P}}_{r,n}$ denotes the set of size $r$ equipartitions of $rn]\!]:=\{1,2,\dots rn\}$. That is, $\{\mathcal{A}_{1},\dots\mathcal{A}_{r}\}\in{\mathcal{P}}_{r,n}$ if $|\mathcal{A}_{i}|=n$ for all $i$ and the $\mathcal{A}_{i}$ form a partition of [[rn]][\![rn]\!]. With that, for any $\mathcal{A}\subset[\![rn]\!]$, we write $\Delta(\mathcal{A})$ as shorthand for the Vandermonde determinant in the $|\mathcal{A}|$ ordered eigenvalue variables with indices drawn from $\mathcal{A}$ (suppressing the explicit dependence on $\lambda_{i},i\in\mathcal{A}$). Finally, $\operatorname{Pf}(M)$ denotes the Pfaffian of $M$.
@@ 1.1 Soft edge asymptotics of $\mathtt{H}_{\beta,n}(r,s)$
When $\gamma=1$ this is the multivariate Stochastic Airy Operator of [2]. In particular, with a Dirichlet boundary condition at the origin, the spectrum of $\mathcal{H}_{\beta}=-\mathcal{H}_{\beta,1}$ is given by the $\operatorname{Airy}_{\beta}$ process, the edge scaling limit of the Gaussian beta ensemble. The largest value of this process (which is minus the ground state eigenvalue of $\mathcal{H}_{\beta}$), has classical Tracy-Widom distribution $TW_{\beta}$ with $\beta=1,2,4$.
@@
and by $\Lambda_{0}<\Lambda_{1}<\cdots$ the Dirichlet eigenvalues of $\mathcal{H}_{\beta,\gamma}$ with the choice $\gamma=\frac{r+s}{r}$ . Then the point process $\{\lambda_{0}^{(n)},\lambda_{1}^{(n)},\dots\}$ converges in law to $\{\Lambda_{0},\Lambda_{1},\dots\}$ as $n\to\infty$.
@@
starting from $(p_{1}(0),\cdots,p_{r}(0))=\{\infty\}^{r}$ and entering $\{p_{1}<\cdots<p_{r}\}$ at $x>0$. Here $(b_{1},\cdots b_{k})$ is a standard real $r$-dimensional Brownian motion; $p_{1}$ can hit -$\infty$ in finite time, whereupon it is placed at $\infty$ and the re-indexed process starts afresh. Then with $\Lambda_{0}<\Lambda_{1}<\cdots$ defined as in Theorem 2, it holds that
@@ 1.2 Block Laguerre ensembles
When $\beta=1$ or $2$ this coincides with that of the law of a sample covariance matrix for $m\geq n$ independent real or complex normal samples in dimension $n$. Along with $\beta$ now taking any positive value, the model behind (7) allows $m$ to be generalized to any real number greater than n1n-1.
We define the distribution $\mathtt{W}_{n,m,\beta}(r,s)$ on nonnegative definite block tridiagonals as follows. Let $\mathbf{L}_{n}$ be an $rn\times rn$ block bidiagonal matrix with independent $r\times r$ diagonal and upper offdiagonal blocks denoted by $\{\mathbf{D}_{i}\}_{i=1,n}$ and $\{\mathbf{O}_{i}\}_{i=1,n-1}$, that are lower and upper triangular matrices, respectively. Distribute these according to square-root Wishart matrices with parameters (r,(r+s)(m+1i))(r,(r+s)(m+1-i)) and (r,(r+s)(ni))(r,(r+s)(n-i)), respectively. Then $\mathtt{W}_{n,,m,\beta}(r,s)$ has the law $\mathbf{L}_{n}\mathbf{L}_{n}^{\dagger}$. Full details are provided in Definition 20.
@@
The joint eigenvalue density of $\mathtt{W}_{n,m,\beta}(r,s)$ for $\beta=1$ or 22 has the form (2) for general $r\geq 2$ and $\beta s=2$ and (3) for $r=2$ and $\beta s=2$ or 44 with an explicitly computable normalizing constant, the only change being that the Gaussian weight $e^{-\frac{\beta}{4}\sum_{i=1}^{rn}\lambda_{i}^{2}}$ is replaced by $\prod_{i=1}^{rn}\lambda_{i}^{\frac{\beta}{2}((r+s)(m-n)+1)-1}e^{-\frac{\beta}{2}\lambda_{i}}$, restricted to ${\mathbb{R}}_{+}^{rn}$.
In terms of asymptotics, we focus on the choice $m=n+a$ for fixed a>1/(r+s)a>-1/(r+s) as $n\rightarrow\infty$ and look at the scaling limit of the smallest eigenvalues, which end up being in the vicinity of the origin. This is the random matrix hard edge, and introduces novel limiting phenomena beyond what we have seen for $\mathtt{H}_{n,\beta}(r,s)$. Note that it may proved along the same lines to Theorem 2 that the suitably centered and scaled largest eigenvalues under $\mathtt{W}_{n,m,\beta}(r,s)$ will converge to those of $\mathcal{H}_{\beta,\gamma}$, for an appropriate $\gamma$, and the same is in fact true for the smallest eigenvalues when $\liminf_{n\to\infty}m/n>1$.
@@
Let $\mathbb{P}$ be the measure on (non-intersecting) paths ${\mathbf{q}}:[0,\infty)\mapsto[-\infty,\infty]^{r}$ defined by
$dq_{i}=\frac{2}{\sqrt{\beta\gamma}}q_{i}db_{i}+\left(\left(\frac{a}{\gamma}+\frac{2}{\beta\gamma}\right)q_{i}-q_{i}^{2}-\lambda e^{-rx}+q_{i}\sum_{j\neq i}\frac{q_{i}+q_{j}}{q_{i}-q_{j}}\right)dx,$ (9)
started from $\{\infty\}^{r}$ with the same ordering and re-indexing conventions upon possible passages to -$\infty$ described in Corollary 3. With $0<\hat{\Lambda}_{0}<\hat{\Lambda}_{1}<\cdots$ defined in Theorem 7 it holds
[…]
@@ 2.1 Block Jacobi matrices
$\int x^{j}d\mu_{[\![r]\!]}=\mathbf{e}_{[\![r]\!]}^{{\dagger}}\mathbf{M}^{j}\mathbf{e}_{[\![r]\!]}=\mathbf{e}_{[\![r]\!]}^{{\dagger}}\mathbf{T}^{j}\mathbf{e}_{[\![r]\!]}.$ (20)
@@ 2.2 Random block matrices
Let $r,n\geq 1$, m>1/rm>-1/r, and $s\geq 0$. Let $\mathbf{L}$ be an $rn\times rn$ block bidiagonal matrix with independent $r\times r$ diagonal and upper offdiagonal blocks denoted by $\{\mathbf{D}_{i}\}_{i=1,n}$ and $\{\mathbf{O}_{i}\}_{i=1,n-1}$ with $\mathbf{D}_{i}^{{\dagger}}\sim\mathtt{SqW}_{\beta}(r,(r+s)(m+1-i))$ and $\mathbf{O}_{i}\sim\mathtt{SqW}_{\beta}(r,(r+s)(n-i))$. We denote the distribution of $\mathbf{W}=\mathbf{L}\mathbf{L}^{{\dagger}}$ by $\mathtt{W}_{\beta,n,m}(r,s)$.
@@ 3.1 Revisiting the Hermite beta ensemble
[…]
@@
[…]
@@ 3.2 Key spectral identity
[…]
@@ 3.3 Proofs of Theorem 1 and 6
$\times\prod_{j=1}^{n-1}(\det\mathbf{O}_{j})^{\beta(r+s)(n-j)+\beta-1}\prod_{j=1}^{n}\prod_{k=1}^{r}(d^{(j)}_{k,k})^{-\beta k}\prod_{j=1}^{n-1}\prod_{k=1}^{r}(o^{(j)}_{k,k})^{-\beta k},$
@@ 3.4 A reduction and a representation
For $\underline{\mathcal{A}}=(\mathcal{A}_{1},\dots,\mathcal{A}_{r})\in{\mathcal{P}}_{r,n}$ let $\sigma_{\underline{\mathcal{A}}}\in S_{rn}$ denote the following permutation of the set [[rn]][\![rn]\!]:
@@
Now when $r=2$ the elements of $(\mathcal{A}_{1},\mathcal{A}_{2})$ of ${\mathcal{P}}_{2,n}$ can be identified just with the subsets $\mathcal{A}\subset[\![2n]\!]$ of cardinality $n$. Using $\mathcal{A}^{\prime}$ for the complement of $\mathcal{A}$ in [[2n]][\![2n]\!] throughout, we further denote:
@@
First, any permutation $\sigma\in S_{rn}$ of [[rn]][\![rn]\!] can be mapped in a one-to-one way to a pair consisting of an element of ${\mathcal{P}}_{r,n}$ and a vector $(\sigma_{1},\dots,\sigma_{r})$ of $r$ permutations of $S_{n}$. For $\sigma\in S_{rn},m\in[\![r]\!]$ let
@@
$\hskip 130.0pt\times\Delta(\mathcal{A})\Delta(\mathcal{A}^{\prime})\Delta(\mathcal{B})\Delta(\mathcal{B}^{\prime}).$
@@
$\hskip 80.0pt\times\sum_{a=0}^{n}g(a,a,2n-a)\sum_{|\mathcal{B}|=n,|\mathcal{A}\cap\mathcal{B}|=a}\operatorname{sgn}(\sigma_{\mathcal{B}})\Delta(\mathcal{B})\Delta(\mathcal{B}^{\prime}).$
@@
To prove (60) in the $1\leq a\leq n-1$ case we use the following identity due to Sylvester [24]. Suppose that $\mathbf{R}_{1}$, $\mathbf{R}_{2}$ are $n\times n$ matrices and $1\leq m\leq n$. Let $I$ be a fixed subset of column indices of $\mathbf{R}_{1}$ with $|I|=m$. Then
@@
Fix $1\leq a\leq n-1$, $\mathcal{A}\subset[\![2n]\!]$ with $|\mathcal{A}|=n$, and $\mathcal{S}\subset\mathcal{A}$ with $|\mathcal{S}|=a$. Let $\mathbf{R}_{1}$ be the $n\times n$ Vandermonde matrix corresponding to $\mathcal{A}$, $\mathbf{R}_{2}$ the Vandermonde matrix corresponding to $\mathcal{A}^{\prime}$, and consider Sylvester’s identity applied for m=nam=n-a where the fixed set of coordinates of $\mathbf{R}_{1}$ correspond to $\widetilde{\mathcal{S}}:=\mathcal{A}\setminus\mathcal{S}$. For each $(\widehat{\mathbf{R}}_{1},\widehat{\mathbf{R}}_{2})$ in the sum, let $\mathcal{B}$ be the set of indices corresponding to the columns of $\widehat{\mathbf{R}}_{1}$, then $|\mathcal{B}|=n,\mathcal{A}\cap\mathcal{B}=\mathcal{S}$. We set $\mathcal{T}=\mathcal{A}^{\prime}\cap\mathcal{B}^{\prime}$ and $\widetilde{\mathcal{T}}=\mathcal{A}^{\prime}\setminus\mathcal{T}$. The left side of (62) is $\Delta(\mathcal{A})\Delta(\mathcal{A}^{\prime})$, while the sum on the right side is
@@
[…]
@@
$\mathcal{B}_{(j_{1},j_{2},j_{3},j_{4})}=\mathcal{A}_{1}^{*}\cap\mathcal{A}_{2}^{*}\cap\mathcal{A}_{3}^{*}\cap\mathcal{A}_{4}^{*},\quad\text{with}\quad\mathcal{A}_{i}^{*}=\begin{cases}\mathcal{A}_{i},&\qquad\text{if }j_{i}=1,\\ \mathcal{A}_{i}^{\prime},&\qquad\text{if }j_{i}=0.\end{cases}$ (67)
@@
In other words, after we take expected value in (66), only those quadruples $(\mathcal{A}_{1},\dots,\mathcal{A}_{4})$ contribute for which intersections of the form $\mathcal{A}_{i_{1}}^{\prime}\cap\mathcal{A}_{i_{2}}\cap\mathcal{A}_{i_{3}}\cap\mathcal{A}_{i_{4}}$ and $\mathcal{A}_{i_{1}}\cap\mathcal{A}_{i_{2}}^{\prime}\cap\mathcal{A}_{i_{3}}^{\prime}\cap\mathcal{A}_{i_{4}}^{\prime}$ are empty (for distinct $i_{1},\dots,i_{4}$). Let $\mathcal{G}_{n}$ denote the set of quadruples $(\mathcal{A}_{1},\dots,\mathcal{A}_{4})$ that satisfy this compatibility condition together with $|\mathcal{A}_{i}|=n$, $\mathcal{A}_{i}\subset[\![2n]\!]$.
@@
[…]
Hence $\mathcal{B}_{(1,0,1,0)}$ and $\mathcal{B}_{3}$ both denote the set $\mathcal{A}_{1}\cap\mathcal{A}_{2}^{\prime}\cap\mathcal{A}_{3}\cap\mathcal{A}_{4}^{\prime}$, and $b_{3}=b_{(1,0,1,0)}$. At the same time, if $(\mathcal{A}_{1},\dots,\mathcal{A}_{4})\in\mathcal{G}_{n}$, then each $\mathcal{A}_{i}$ and $\mathcal{A}_{i}^{\prime}$ can be written as the disjoint union of four of the sets $\mathcal{B}_{j},1\leq j\leq 8$. For example, $\mathcal{A}_{1}=\mathcal{B}_{1}\cup\mathcal{B}_{2}\cup\mathcal{B}_{3}\cup\mathcal{B}_{4}$. From $|\mathcal{A}_{i}|=|\mathcal{A}_{i}^{\prime}|$, $1\leq i\leq 4$, we get four equations with the sum of four terms on both sides. For instance, $|\mathcal{A}_{1}|=|\mathcal{A}_{1}^{\prime}|$ yields
@@
Next, for a particular $(\mathcal{A}_{1},\dots,\mathcal{A}_{4})\in\mathcal{G}_{n}$ consider the following “special” ordering of the elements of [[2n]][\![2n]\!]: $x\prec y$ if $x\in\mathcal{B}_{i},y\in\mathcal{B}_{j}$ with $i<j$ or if $x<y$ with $x,y\in\mathcal{B}_{i}$. Denote by $\widetilde{\Delta}(\mathcal{A})$ the Vandermonde determinant corresponding to the elements $\lambda_{i},i\in\mathcal{A}$ using our special ordering, and let $\sigma_{i}$ be the permutation that maps $1,3,\dots,2n-1$ and $2,4,\dots,2n$ into the elements of $\mathcal{A}_{i}$ and $\mathcal{A}_{i}^{\prime}$ according to the special order. These permutations can be visualized using the following table:
@@
Here $a_{1},a_{2}$ are indices from the $\mathcal{B}$-partition” of $\mathcal{A}_{i}$ and $b_{1},b_{2}$ are indices from the the $\mathcal{B}$-partition” of $\mathcal{A}_{i}^{\prime}$. Multiplying these equations for $1\leq i\leq 4$ and some bookkeeping yields